OpenStudy (anonymous):

Are you able to simplfy this one? -14x^3/x^3-5x^4

OpenStudy (anonymous):

factor out x^3 in numerator and denom

OpenStudy (mertsj):

\[\frac{x^3(-14)}{x^3(1-5x)}=\frac{-14}{1-5x}=\frac{14}{5x-1}\]